Determine the amount of an investment if $500 is invested At an interest rate of 4.25% compounded quarterly for 12 years.
Answer by stanbon(75887)

Determine the amount of an investment if $500 is invested At an interest rate of 4.25% compounded quarterly for 12 years.
-----
A(t) = P(1+(r/n))^(nt)
-----
A(12) = 500(1+(0.0425/4))^(4*12)
-----
A(12) = 500(1.010525)^(48)
-----
A(12) = $830.41
===================
Cheers,
Stan H.
===================